# Cluster-cyclic condition of skew-symmetrizable matrices of rank 3 via the Markov constant

### v1

@@ 1.1. Background
where $x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{R}$ satisfy $xyz=x^{\prime}y^{\prime}z^{\prime}$. In the usual cluster algebra theory, we often suppose that all entries of $B$ are integral, since entries of $B$ appear in the exponent of rational fractions. However, since we do not focus on cluster variables, we consider real skew-symmetrizable matrices. When we focus on only matrix mutations $\mu_{k}(B)$, we can naturally extend the definition, and their properties are helpful for the study of cluster algebras. This generalization for real entries have already appeared in [BBH11, FT19, FT23] for the skew-symmetric case.
@@
In the skew-symmetric case, this problem was solved in [BBH11, FT19]. In [BBH11], they showed it for the integer skew-symmetric case, and in [FT19], they generalized it for real skew-symmetric case by using the geometric aspect of mutations. They showed that the Markov constants, which was introduced in [BBH11] for the cluster algebra theory, play important roles in this problem. For the skew-symmetrizable case, the Markov constants $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|$, which are a generalization of the one for skew-symmetric matrices, also play important roles in this problem.
@@ 1.2. Method
However, by considering skew-symetrizable matrices, we need some reconstructions. One of the main reasons is that the surjectivity of Markov constants. (See Theorem 10.12, Example 10.18.) In [BBH11], it was shown that there are no integer skew-symmetric and cluster-cyclic matrices with $40<C(B)<$ 4. Thanks to this property, some arguments became clear. However, by considering skew-symmetrizable matrices, there are both cluster-cyclic and cluster-acyclic matrices $B$ with $40<C(B)<$ 4.
@@ 1.3. Main results
First, we mention the identification of cluster-cyclicity. It is a natural generalization of [BBH11] for the skew-symmetrizable and real case.
Theorem 1.1 (cf. [BBH11, Thm. 1.2 (1),(2)]).
Let $B$ be a skew-symmetrizable and cyclic matrix of the form (1). Then, the following two conditions are equivalent.
@@
For some thechnical reasons, it is better to consider transformations $\gamma_{k}=-\mu_{k}$ than the usual mutations $\mu_{k}$ for any $k=1,2,3$. Let $\Gamma$ be a group generated by $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$. When $B$ of the form (1) satisfies $x,y,z,x^{\prime},y^{\prime},z^{\prime}>0$, $B$ is called positive-cyclic. By considering $\gamma_{k}$, if $B$ is positive and cluster-cyclic, then every element in this orbit $\Gamma(B)$ is also so. (See Proposition 4.1.)
Next result is about a fundamental domain for this action $\Gamma$ of integer skew-symmetrizable matrices. The following partial order $\leq$ on skew-symmetrizable and positive-cyclic matrices is useful in this study.
@@
(3) $\left\{B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\\ \end{matrix}\right)\mid\begin{aligned} &\textup{$B$ is skew-symmetrizable},\\ &xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime},\\ &x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{Z}_{>0}\end{aligned}\right\}$
@@
Nemely, for each integer skew-symmetrizable, positive, and cluster-cyclic matrix $B$, there is the minimum element of $\Gamma(B)$ and it is an element in $\tilde{F}$. Unfortunately, this result cannnot generalize for the real case. (See Theorem 9.1.)
These two results are natural generalizations of the skew-symmetric case. On the other hand, the following result is the most important difference between the skew-symmetrizable case and the skew-symmetric case.
@@
The Markov constant $C(B)$ can take any integer less than or equal to $4$ by choosing an integer skew-symmetrizable and cluster-cyclic matrix $B$.
Because of this property (in particular, the fact that there is a cluster-cyclic matrix $B$ with $40<C(B)<$ 4), we need to reconstruct a proof in [BBH11] even though we focus on the integer skew-symmetrizable matrices (see Example 10.18).
@@ 1.4. Structure of the paper
In section 4, we introduce more simple notations, which is similar to [BBH11, FT19]. In this notations, Theorem 1.1 is decomposed to Lemma 4.7, Lemma 4.8, and Lemma 4.9.
@@
In section 6, we calculate some examples and prove Lemma 4.8. Most contents and ideas were in [FT19].
@@
In section 9 and 10, we derive some new results. In section 9, we consider the case which does not occur when we consider integer skew-symmetric matrces. The related topic is in [FT19]. In section 10, we deal with the integer skew-symmetrizable case. In particular, we focus on the difference between the result of the skew-symmetric case in [BBH11] and the one of the skew-symmetrizable case.
@@ 2.1. Skew-symmetrizable matrix and mutation
In this paper, we will focus on skew-symmetrizable matrices of rank $3$ with real entries. Let $\mathrm{M}_{3}(\mathbb{R})$ be the set of all $3\times 3$ real matrices. A real matrix $B=(b_{ij})\in\mathrm{M}_{3}(\mathbb{R})$ is called skew-symmetrizable if there exists a diagonal matrix $D=\mathrm{diag}(d_{1},d_{2},d_{3})$ with positive real numbers $d_{1},d_{2},d_{3}>0$ such that $DB$ is skew-symmetric. This $D$ is said to be a skew-symmetrizer of $B$.
@@
(a) $B$ is sign-skew-symmetric, namely, $b_{ij}$ and $-b_{ji}$ have the same sign for any $i,j$. (In particular, every diagonal entry is 00.)
@@ 3. Markov constant
For any skew-symmetrizable matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ of the form (13), we define the Markov constant $C(B)$ of $B$ as follows:
@@
Similarly, we define $C^{+}(B)$ and $C^{-}(B)$ as
(15) $C^{\pm}(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-(\pm xyz).$
These three numbers have the following relations.
@@
(a) We have $C(B)=C(-B)$ and $C^{+}(B)=C^{-}(-B)$.
(b) If $B$ is positive-cyclic (resp. negative-cyclic), then we have $C(B)=C^{+}(B)$ (resp. $C(B)=C^{-}(B)$).
We may obtain these properties by considering a sign of $xyz$.
For any skew-symmetrizable matrix $B$, since $xyz=x^{\prime}y^{\prime}z^{\prime}=pqr$ as shown in (10), the Markov constant $C(B)$ may also be expressed as follows:
(16) $C(B)$ $=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|$
$=xx^{\prime}+yy^{\prime}+zz^{\prime}-|x^{\prime}y^{\prime}z^{\prime}|$
$=p^{2}+q^{2}+r^{2}-|pqr|=C(\mathrm{Sk}(B)).$
The following proposition is fundamental for later.
Proposition 3.3.
Let $B$ be a skew-symmetrizable matrix. Then, the following two statements hold.
(a) We have $C^{\pm}(B)=C^{\pm}(\mathrm{Sk}(B))$ and $C(B)=C(\mathrm{Sk}(B))$.
(b) If $B$ is positive-cyclic (resp. negative-cyclic), we have $C^{-}(\mu_{k}(B))=C^{+}(B)$ (resp. $C^{+}(\mu_{k}(B))=C^{-}(B)$) for any $k=1,2,3$. In particular, we have $C^{+}(-\mu_{k}(B))=C^{+}(B)$ if $B$ is positive-cyclic.
@@
Let $B$ be expressed as in (13). The statement (a) has already been shown in (16). We show (b) in the case of $k=1$. Let $\varepsilon$ be $+$ and -- if $B$ is positive-cyclic and negative-cyclic, respectively. Since $C^{\varepsilon}(B)=C^{\varepsilon}(\mathrm{Sk}(B))$ and $C^{-\varepsilon}(\mu_{1}(B))=C^{-\varepsilon}(\mathrm{Sk}(\mu_{1}(B)))=C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$ by (a) and Proposition 2.6 (b), it suffices to show that $C^{\varepsilon}(\mathrm{Sk}(B))=C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$. We have
(17) $\mu_{1}(\mathrm{Sk}(B))=\left(\begin{matrix}0&r&-q\\ -r&0&-p+\varepsilon qr\\ q&p-\varepsilon qr&0\end{matrix}\right).$
Thus, we obtain
(18) $C^{-\varepsilon}(\mu_{1}(\mathrm{Sk}(B)))$ $=(p-\varepsilon qr)^{2}+q^{2}+r^{2}-(-\varepsilon)(p-\varepsilon qr)(-q)(-r)$
$=p^{2}+q^{2}+r^{2}-\varepsilon pqr=C^{\varepsilon}(\mathrm{Sk}(B)).$
4. Reduction of Theorem 1.1 to the skew-symmetric case
@@
Let $B\in\mathrm{M}_{3}(\mathbb{R})$ be a skew-symmetrizable matrix. Let $t=(t_{1},t_{2},\dots,t_{l})$ be a finite sequence of indices $1,2,3$ such that $t_{i}\neq t_{i+1}$ for any $i=1,2,\dots,l-1$. The length $|t|$ is defined by the number of indices; in this case, the length of $t$ is $l$. We write $\mu_{t}(B)=\mu_{t_{l}}\cdots\mu_{t_{2}}\mu_{t_{1}}(B)$. A collection $\Sigma(B)=\{\mu_{t}(B)\}_{t}$ indexed by all of such $t$ is called the matrix-pattern of $B$. In convention, we allow $t=\emptyset$ and let $\mu_{\emptyset}=\mathrm{id}$. We often fix one element of $\Sigma(B)$ and we call it an initial matrix of $\Sigma(B)$. Now, we may see $B$ as an initial marix of $\Sigma(B)$. Then the distance (from $B$) to $\mu_{t}(B)\in\Sigma(B)$ is defined by the length of $t$.
Let
be a skew-symmetrizable and cyclic matrix. Recall that $B$ is said to be positive-cyclic (resp. negative-cyclic) if all of $x,y,z,x^{\prime},y^{\prime},z^{\prime}$ are positive (resp. negative). Later, we often consider positive-cyclic matrices. If we consider a negative-cyclic matrix $B$, apply the results for $-B$. Our interest is whether a given positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{Z})$ is cluster-cyclic or not. In this case, it is enough to consider the following maps $\gamma_{k}$ instead of $\mu_{k}$.
(20) $\gamma_{1}(B)$ $=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-(yz-x^{\prime})\\ -y^{\prime}&y^{\prime}z^{\prime}-x&0\end{matrix}\right),$
$\gamma_{2}(B)$ $=\left(\begin{matrix}0&-z^{\prime}&z^{\prime}x^{\prime}-y\\ z&0&-x^{\prime}\\ -(zx-y^{\prime})&x&0\end{matrix}\right),$
$\gamma_{3}(B)$ $=\left(\begin{matrix}0&-(xy-z^{\prime})&y\\ x^{\prime}y^{\prime}-z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),$
where the domain of these functions consists of all skew-symmetrizable matrices (not necessarily cyclic).
We may verify that $\gamma_{k}\gamma_{k}(B)=B$. Define $\Gamma(B)=\{\gamma_{t}(B)\}_{t}$ in the same way as $\Sigma(B)$, but we replace $\mu_{k}$ with $\gamma_{k}$. In this case, we have the following proposition.
Proposition 4.1.
Let $B$ be a skew-symmetrizable and positive-cyclic matrix.
(a) We have $\gamma_{k}(B)=-\mu_{k}(B)$ for any $k=1,2,3$.
(b) The following two conditions are equivalent.
  • (1)
    $B$ is cluster-cyclic.
  • (2)
    Every matrix of $\Gamma(B)$ is positive-cyclic.
Proof.
We may show (a) by the direct calculation. By considering (a), we may show (b) as follows: Note that $\mu_{k}(-B)=-\mu_{k}(B)$. For a sequence of indices $t=(t_{1},\dots,t_{l})$ and $t^{\prime}=(t_{1},\dots,t_{l},t_{l+1})$, if $\gamma_{t}(B)=(-1)^{|t|}\mu_{t}(B)$ is cluster-cyclic, we have $\gamma_{t^{\prime}}(B)=(-1)^{|t^{\prime}|}\mu_{t^{\prime}}(B)$. This relation is true when $\mu_{t}(B)$ is cyclic. (we can disregard the condition of whether $\mu_{t^{\prime}}(B)$ is cyclic or not.) Thus, if there is an acyclic-matrix $\mu_{t^{\prime}}(B)$ and $|t^{\prime}|$ is the smallest among such $t^{\prime}$, then it holds that $\gamma_{t^{\prime}}(B)=(-1)^{|t^{\prime}|}\mu_{t^{\prime}}(B)$ and it is acyclic. ∎
Later, we write a skew-symmetrizable matrix by a following tuplet:
@@
The maps $\gamma_{k}$ may be expressed as
@@
is cluster-cyclic or not. We may see this matrix as a triplet of real numbers $(p,q,r)\in\mathbb{R}^{3}$ and the maps $\gamma_{k}$ become
@@
+=$\mathcal{M}^{+}=\left\{\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in\mathcal{M}\mid x,y,z,x^{\prime},y^{\prime},z^{\prime}>0\right\},$
$\mathcal{S}=\{(p,q,r)\mid p,q,r\in\mathbb{R}\}=\mathbb{R}^{3},$
$\mathcal{S}^{+}=\{(p,q,r)\in\mathcal{S}\mid p,q,r>0\}.$
@@
The maps $\gamma_{k}$ satisfy $\gamma_{k}^{2}=\mathrm{id}$, in particular, $\gamma_{k}$ is a bijection on $\mathcal{S}$.
@@
For any $M\in\mathcal{M}$, let $\Gamma(M)$ be the collection of $\gamma_{t}(M)$ indexed by all of finite sequences $t=(t_{1},\dots,t_{k})$ of indices $1,2,3$ with $t_{i}\neq t_{i+1}$, where $\gamma_{t}=\gamma_{t_{k}}\cdots\gamma_{t_{1}}$. For any positive element $M\in\mathcal{M}^{+}$, if every element in $\Gamma(M)$ is positive, $M$ is said to be cluster-positive. Similarly, we define the cluster-positivity of $S\in\mathcal{S}^{+}$.
As we said, every positive element $M\in\mathcal{M}^{+}$ and $S\in\mathcal{S}^{+}$ corresponds to a positive-cyclic matrices. Here, we define a map $\mathrm{Sk}:\mathcal{M}\to\mathcal{S}$ as
@@
where $\varepsilon_{a}\in\{0,\pm 1\}$ is the sign of $a\in\mathbb{R}$. Obviously, if $M\in\mathcal{M}$ corresponds to a positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{R})$, then $\mathrm{Sk}(M)$ corresponds to its double-sided skew-symmetrized form $\mathrm{Sk}(B)$. Moreover, $M$ is cluster-positive if and only if $S=\mathrm{Sk}(M)$ is cluster-positive.
@@
Proof is similar to the one of Proposition 3.3 (b), in particular, (18). Thanks to this result, the Markov constant does not change under the transformations of $\gamma_{k}$.
@@
(Proposition 4.1) $B$ is cluster-cyclic if and only if $S$ is cluster-cyclic.
@@
Let $S=(p,q,r)\in\mathcal{S}^{+}$. Then, $S$ is cluster-cyclic if and only if both $p,q,r\geq 2$ and $C(S)\leq 4$ hold.
@@
Let $B$ be a skew-symmetrizable and cyclic matrix. If $B$ is positive-cyclic, the claim holds by using the correspondence of $\mathrm{Sk}(B)$ and $S$. When $B$ is negative-cyclic, then $-B$ is positive-cyclic. Thus, by applying the result of Proposition 4.6 for $-B$, we can show Theorem 1.1. ∎
@@
Here, we show Lemma 4.7. The essential part is as follows.
@@
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy $p,q,r\geq 2$ and $C(S)\leq 4$ and set $S^{\prime}=(p^{\prime},q^{\prime},r^{\prime})=\gamma_{k}(S)$. By definition of $\gamma_{k}$, all entries other than $k$th one do not change. In particular, they are larger or equal to $2$. By Proposition 4.5 (c), $C(S^{\prime})=C(S)\leq 4$ holds. Thus, the $k$th entry of $S^{\prime}$ is also larger or equal to $2$ by Lemma 5.1. By repeating this argument, every entry of elements in $\Gamma(S)$ is larger or equal to $2$ and, in particular, positive. Thus, the claim holds. ∎
6. $1,2$-orbits and proof of Lemma 4.8
This contents was introduced in [FT19, Lem. 3.3].
@@
Let $S\in\mathcal{S}^{+}$. We define a collection $\Gamma_{1,2}(S)=\{\gamma_{t}(S)\}_{t}$ indexed by a sequence $t=(t_{1},\dots,t_{l})$ of $1,2$ such that $t_{i}\neq t_{i+1}$ for any $i=1,\dots,l-1$.
@@
We may check it by the induction on $n$. By using this property, entries of $\Gamma_{1,2}(S)$ are closely related to the Chebychev polynomial. The following properties are well known for the Chebyshev polynomials.
@@
For the reader’s convenience, we give a proof following [FT19]. When $20<r<$ 2, we may express $r=2\cos\theta$ for some $\theta\in\mathbb{R}$ with $0<\theta<\frac{\pi}{2}$. By Proposition 6.3, $f_{n}(S)=qu_{n}(r)-pu_{n-1}(r)$ is a entry of an element in $\Gamma(S)$. We show that there is $n\in\mathbb{Z}_{\geq 0}$ such that $f_{n}(S)<0$. Note that $u_{n}(r)=U_{n}(\cos\theta)=\frac{\sin(n+1)\theta}{\sin\theta}$ by (38) and Proposition 6.4. By using this relation, the inequality $f_{n}(S)<0$ is equivalent to $q\sin(n+1)\theta<p\sin n\theta$. Here we use $\sin\theta>0$ since $0<\theta<\frac{\pi}{2}$. We can find $n\geq 1$ such that $\sin(n+1)\theta\leq 0$ while every $k=1,\dots,n$ satifies $\sin k\theta>0$. This $n$ satisfies $q\sin(n+1)\theta<p\sin n\theta$. ∎
@@
Let $S=(p,q,r)\in\mathcal{S}^{+}$ with $p,q,r\geq 2$ and set $r=2\cosh\theta$ for some $\theta\in\mathbb{R}_{\geq 0}$. Let $p_{n}$ and $q_{n}$ be in (41). If $p_{n}$ or $q_{n}$ converges to some real number, then we have $qe^{\theta}-p=0$. Moreover,
@@
(b) if $r>2$, then both $p_{n}$ and $q_{n}$ converge to 00.

### v2

@@ 1.1. Background
where $x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{R}$ satisfy $xyz=x^{\prime}y^{\prime}z^{\prime}$. In the usual cluster algebra theory, we often suppose that all entries of $B$ are integral, since entries of $B$ appear in the exponent of rational fractions. However, since we do not focus on cluster variables, we may consider real skew-symmetrizable matrices. When we focus on only matrix mutations $\mu_{k}(B)$, we can naturally extend the definition, and their properties are helpful for the study of cluster algebras. This generalization for real entries have already appeared in [BBH11, FT19, FT23] for the skew-symmetric case.
@@
1.2. Previous works
In the skew-symmetric case, this problem was solved in [BBH11, FT19]. In [BBH11], they showed it for the integer skew-symmetric case, and in [FT19], they generalized it for real skew-symmetric case by using the geometric aspect of mutations. They showed that the Markov constants, which was introduced in [BBH11] for the cluster algebra theory, play important roles in this problem. For the skew-symmetrizable case, the Markov constants $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|$, which are a generalization of the one for skew-symmetric matrices, also play important roles in this problem. For the skew-symmetrizable case, this problem was solved by Seven [Sev12] by using a quasi-Cartan companion.
@@ 1.2. Previous works
Remark 1.1.
After the author submitted this paper, I was informed by Ahmet Seven of the related works in [Sev12, Sev13]. By using results in [Sev12], Theorem 1.2 can be shown. A related result to Theorem 9.7 appeared in [Sev13]. However, to obtain other results, we necessitate a different proof that draws upon certain facts from the proof of Theorem 1.2. So, we give an alternative proof of Theorem 1.2 and Theorem 9.7 in this paper.
@@ 1.3. Method
However, by considering skew-symetrizable matrices, we need some reconstructions. One of the main reasons is that the surjectivity of Markov constants. (See Theorem 9.12, Example 9.18.) In [BBH11], it was shown that there are no integer skew-symmetric and cluster-cyclic matrices with $0<C(B)<4$. This property simplifies certain arguments compared to the skew-symmetrizable case. However, due to this difference, we require some modifications to its proof.
@@ 1.4. Main results
First, we mention the identification of cluster-cyclicity. It is a natural generalization of [BBH11] for the skew-symmetrizable and real case. The following result has already obtained in [Sev12] essentially by using a quasi-Cartan companion.
Theorem 1.2 (cf. [BBH11, Thm. 1.2 (1),(2)], [Sev12, Thm. 2.6]).
Let $B$ be a skew-symmetrizable and cyclic matrix of the form (1). Define its Markov constant
(2) $C(B)=xx^{\prime}+yy^{\prime}+zz^{\prime}-|xyz|.$
Then, the following two conditions are equivalent.
@@
For some thechnical reasons, it is better to consider transformations $\gamma_{k}=-\mu_{k}$ than the usual mutations $\mu_{k}$ for any $k=1,2,3$. Let $\Gamma$ be a group generated by $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$. When $B$ of the form (1) satisfies $x,y,z,x^{\prime},y^{\prime},z^{\prime}>0$, $B$ is called positive-cyclic. By considering $\gamma_{k}$, if $B$ is positive and cluster-cyclic, then every element in this orbit $\Gamma(B)$ is also so.
Next result is to give a fundamental domain for this action $\Gamma$ of integer skew-symmetrizable matrices. The following partial order $\leq$ on skew-symmetrizable and positive-cyclic matrices is useful in this study.
@@
(4) $\left\{\left.B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\\ \end{matrix}\right)\ \right|\ \begin{aligned} &\textup{$B$ is skew-symmetrizable},\\ &xyz\geq 2xx^{\prime},2yy^{\prime},2zz^{\prime},\\ &x,y,z,x^{\prime},y^{\prime},z^{\prime}\in\mathbb{Z}_{>0}\end{aligned}\right\}$
@@
Namely, for each integer skew-symmetrizable, positive, and cluster-cyclic matrix $B$, there is the minimum element of $\Gamma(B)$, and $\tilde{F}$ is the set of all such minimum elements. Unfortunately, this result cannnot generalize for the real case. (See Theorem 10.3.)
These two results are natural generalizations of the skew-symmetric case. On the other hand, the following result is the biggest difference between the skew-symmetrizable case and the skew-symmetric case.
@@
The Markov constant $C(B)$ can take any integer less than or equal to 444 by choosing an integer skew-symmetrizable and cluster-cyclic matrix $B$.
[…]
@@ 1.5. Structure of the paper
In section 4, we introduce more simple notations, which is similar to [BBH11, FT19]. In this notations, we may decompose Theorem 1.2 into Lemma 4.7, Lemma 4.8, and Lemma 4.9.
@@
In section 6, we show that some mutation can be expressed by the Chebyshev polynomials of the second kind. Based on this, we may prove Lemma 4.8. These ideas have appeared in [FT19].
@@
In section 9 and 10, we derive some new results. In section 9, we deal with the integer skew-symmetrizable case. In particular, we focus on the difference between the result of the skew-symmetric case in [BBH11] and the one of the skew-symmetrizable case. In section 10, we consider the case which does not occur when we consider integer skew-symmetric matrces.
@@ Acknowledgement.
I thank Ahmet Seven for noticing the important related works.
@@ 2.1. Skew-symmetrizable matrix and mutation
In this paper, we will focus on skew-symmetrizable matrices of rank 333 with real entries. Let $\mathrm{M}_{3}(\mathbb{R})$ be the set of all $\times 3$ 3 real matrices. A real matrix $B=(b_{ij})\in\mathrm{M}_{3}(\mathbb{R})$ is called skew-symmetrizable if there exists a diagonal matrix $D=\mathrm{diag}(d_{1},d_{2},d_{3})$ with positive real numbers $d_{1},d_{2},d_{3}>0$ such that $DB$ is skew-symmetric. This $D$ is said to be a skew-symmetrizer of $B$.
@@
(a) $B$ is sign-skew-symmetric, namely, $b_{ij}$ and $-b_{ji}$ have the same sign for any $i,j$. (In particular, every diagonal entry is 0.)
@@ 3. Markov constant
For any cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{R})$ of the form (14), define the Markov constant
@@
The following properties are fundamental.
@@
(a) We have $C(\mathrm{Sk}(B))=C(B)$.
(b) For any $k=1,2,3$, we have $C(\mu_{k}(B))=C(B)$.
@@
The statement (a) follows from (11). We show (b). Set
(16) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),\ \mu_{k}(B)=\left(\begin{matrix}0&w^{\prime}&-v\\ -w&0&u^{\prime}\\ v^{\prime}&-u&0\end{matrix}\right).$
Suppose that $B$ is positive-cyclic. Regardless $B$ and $\mu_{k}(B)$ is cyclic or not, we have $\varepsilon_{\mu_{k}(B)}|uvw|=uvw$. Thus, we have
(17) $C(\mu_{k}(B))=uu^{\prime}+vv^{\prime}+ww^{\prime}-uvw.$
We can check $C(\mu_{k}(B))=C(B)$ by a direct calculation. For example, if $k=1$, we have
(18) $C(\mu_{1}(B))$ $=(y^{\prime}z^{\prime}-x)(yz-x^{\prime})+yy^{\prime}+zz^{\prime}-(y^{\prime}z^{\prime}-x)yz$
$=xx^{\prime}+yy^{\prime}+zz^{\prime}-x^{\prime}y^{\prime}z^{\prime}=C(B).$
Suppose that $B$ is acyclic. For the simplicity, we assume $x,x^{\prime}\leq 0$ and $y,y^{\prime},z,z^{\prime}>0$. Then, regardless of the cyclicity of $\mu_{k}(B)$, we have $\varepsilon_{\mu_{k}(B)}|uvw|=-uvw$ holds. By a similar calculation of (18), we may show $C(\mu_{k}(B))=C(B)$. ∎
4. Notations
@@
Our purpose is to identify whether a given skew-symmetrizable matrix $B$ is cluster-cyclic or not. In this case, $\gamma_{k}=-\mu_{k}$ is more convenient than the usual mutation $\mu_{k}$. Note that $-\mu_{k}(B)=\mu_{k}(-B)$ for any skew-symmetrizable matrix $B$. Thus, $B$ is cluster-cyclic if and only if every $\gamma_{i_{s}}\cdots\gamma_{i_{1}}(B)=(-1)^{s}\mu_{i_{s}}\cdots\mu_{i_{1}}(B)$ is cyclic.
Moreover, for any skew-symmetrizable matrix
(19) $B=\left(\begin{matrix}0&-z^{\prime}&y\\ z&0&-x^{\prime}\\ -y^{\prime}&x&0\end{matrix}\right),$
we will write it as the following tuplet:
@@
Moreover, $B$ is cyclic if and only if all entries of $M$ have the same sign. If $B$ is cyclic, the maps $\gamma_{k}$ may be expressed as
@@
is cluster-cyclic or not. We may see this matrix as a triplet of real numbers $(p,q,r)\in\mathbb{R}^{3}$. If $\mathrm{Sk}(B)$ is cyclic, the maps $\gamma_{k}$ may be expressed as
@@
Our interest is for cyclic matrices, and we do not treat acyclic matrixes. So, we see (22) and (24) as the definition of $\gamma_{k}$.
@@
$\mathcal{M}^{+}$ =$=\left\{\left.\left(\begin{matrix}x&y&z\\ x^{\prime}&y^{\prime}&z^{\prime}\end{matrix}\right)\in\mathcal{M}\ \right|\ x,y,z,x^{\prime},y^{\prime},z^{\prime}>0\right\},$
$\mathcal{S}$ =$=\{(p,q,r)\mid p,q,r\in\mathbb{R}\}=\mathbb{R}^{3},$
$\mathcal{S}^{+}$ =$=\{(p,q,r)\in\mathcal{S}\mid p,q,r>0\}.$
@@
For any finite sequence $t=(t_{1},\dots,t_{k})$ of indices 1,2,31,2,31 , 2 , 3, set $\gamma_{t}=\gamma_{t_{k}}\cdots\gamma_{t_{1}}$. For any $M\in\mathcal{M}$, let $\Gamma(M)$ be the collection of $\gamma_{t}(M)$ indexed by all of finite sequences $t=(t_{1},\dots,t_{k})$ of indices 1,2,31,2,31 , 2 , 3 with $t_{i}\neq t_{i+1}$, where $\gamma_{t}=\gamma_{t_{k}}\cdots\gamma_{t_{1}}$. For any positive element $M\in\mathcal{M}^{+}$, if every element in $\Gamma(M)$ is positive, $M$ is said to be cluster-positive. Similarly, we define the cluster-positivity of $S\in\mathcal{S}^{+}$.
Here, we define a map $\mathrm{Sk}:\mathcal{M}\to\mathcal{S}$ as
@@
where $\varepsilon_{a}\in\{0,\pm 1\}$ is the sign of $a\in\mathbb{R}$. If $M\in\mathcal{M}$ corresponds to a positive-cyclic matrix $B\in\mathrm{M}_{3}(\mathbb{R})$, then $\mathrm{Sk}(M)$ corresponds to its double-sided skew-symmetrized form $\mathrm{Sk}(B)$. Moreover, $M$ is cluster-positive if and only if $S=\mathrm{Sk}(M)$ is cluster-positive.
@@
(Proposition 2.6) $B$ is cluster-cyclic if and only if $S$ is cluster-positive.
@@
Let $S=(p,q,r)\in\mathcal{S}^{+}$. Then, $S$ is cluster-positive if and only if both $p,q,r\geq 2$ and $C(S)\leq 4$ hold.
Remark 4.6.
This result has essentially generalized for skew-symmetrizable case by [Sev12]. In this paper, the result is expressed by the quasi-Cartan companion instead of the Markov constant.
@@
[…]
@@
Here, we show Lemma 4.7. The essential part is the following.
@@
Let $S=(p,q,r)\in\mathcal{S}^{+}$ satisfy $p,q,r\geq 2$ and $C(S)\leq 4$ and set $S^{\prime}=(p^{\prime},q^{\prime},r^{\prime})=\gamma_{k}(S)$. By definition of $\gamma_{k}$, all entries other than $k$th one do not change. In particular, they are larger or equal to 222. By Proposition 4.4 (c), $C(S^{\prime})=C(S)\leq 4$ holds. Thus, the $k$th entry of $S^{\prime}$ is also larger or equal to 222 by Lemma 5.1. By repeating this argument, every entry of elements in $\Gamma(S)$ is larger or equal to 222 and, in particular, positive. Thus, the claim holds. ∎
6. 1,21,21 , 2-orbits and proof of Lemma 4.8
The contents in this section have already appeared in [FT19, Lem. 3.3].
@@
Let $S\in\mathcal{S}^{+}$. We define a collection $\Gamma_{1,2}(S)=\{\gamma_{t}(S)\}_{t}$ indexed by a sequence $t=(t_{1},\dots,t_{l})$ of 1,21,21 , 2 such that $t_{i}\neq t_{i+1}$ for any $i=1,\dots,l-1$.
@@
We may check it by the induction on $n$. By using this property, entries of $\Gamma_{1,2}(S)$ are closely related to the Chebyshev polynomials. The following properties are well known.
@@
For the reader’s convenience, we give a proof following [FT19]. When $0<r<2$, we may express $r=2\cos\theta$ for some $\theta\in\mathbb{R}$ with $0<\theta<\frac{\pi}{2}$. By Proposition 6.3, $f_{n}(S)=qu_{n}(r)-pu_{n-1}(r)$ is a entry of an element in $\Gamma(S)$. We show that there is $n\in\mathbb{Z}_{\geq 0}$ such that $f_{n}(S)<0$. Note that $u_{n}(r)=U_{n}(\cos\theta)=\frac{\sin(n+1)\theta}{\sin\theta}$ by (37) and Proposition 6.4. By using this relation, the inequality $f_{n}(S)<0$ is equivalent to $q\sin(n+1)\theta<p\sin n\theta$. Here we use $\sin\theta>0$ since $0<\theta<\frac{\pi}{2}$. We can find $n\geq 1$ such that $\sin(n+1)\theta\leq 0$ while every $k=1,\dots,n$ satifies $\sin k\theta>0$. This $n$ satisfies $q\sin(n+1)\theta<p\sin n\theta$. ∎
@@
Let $S=(p,q,r)\in\mathcal{S}^{+}$ with $p,q,r\geq 2$ and set $r=2\cosh\theta$ for some $\theta\in\mathbb{R}_{\geq 0}$. Let $p_{n}$ and $q_{n}$ be in (40). If $p_{n}$ or $q_{n}$ converges to some real numbers, then we have $qe^{\theta}-p=0$. Moreover,
@@
(b) if $r>2$, then both $p_{n}$ and $q_{n}$ converge to 0.